Circular No. 8430

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004fv IN NGC 6492 L. A. G. Monard, Pretoria, S. Africa, reports his discovery of an ap-

parent supernova (mag 14.8 ± 0.2) on unfiltered CCD images obtained on Nov. 4.750 and 6.719 UT with a 0.30-m reflector. The new object is located at α = 18^{h}02^{m}46^{s}.38, δ = −66^{o}25′54′′.7 (equinox 2000.0), which is 12′′ west and 5′′ south of the nucleus of NGC 6492. Nothing was visible at this location on Monard’s image taken on Oct. 16.8 (limiting red mag 18.0) or on a Digitized Sky Survey image (limiting red mag 20.5).

SUPERNOVA 2004ao IN UGC 10862 G. Gomez, Instituto de Astrofisica de Canarias (IAC); R. Lopez, Uni-

versidad de Barcelona; and J. A. Acosta-Pulido, and the LIRIS Team, IAC, report that an infrared image (yielding magnitude J ∼ 16.6) and a spectrogram (range 0.89–1.53 μm; resolution ∼ 700) were obtained of SN 2004ao (cf. IAUC 8299, 8304) on June 8.1 UT with the William Herschel Telescope (+ LIRIS). Preliminary reduction of the spectrum shows absorption and emission features superimposed on a nearly flat continuum, indicating that the supernova had then reached the nebular phase. The spectrum shows a P-Cyg feature at 1.08 μm (absorption at 1.0430 μm and emission at 1.0865 μm, at the rest wavelength of the host galaxy); this feature is identified as He I 1.0830-μm in other core-collapse supernovae (e.g., SN 1999ex, Hamuy et al. 2002, A.J. 124, 417; SN 1998bw, Patat et al. 2001, Ap.J. 555, 900). Adopting the NED host-galaxy recession velocity of 1691 km/s, the expansion velocity derived from the minimum of the He I line is 11000 km/s. Broad emissions at 0.924, 1.089, 1.1295, 1.130, and 1.190 μm and absorptions at 0.9095, 0.979 and 1.0035 μm (all at the host-galaxy rest wavelength), are also detected. Tentative identifications include the 1.1295-μm emission as O I 1.129-μm (e.g., SN 1998bw at phase +51 days, Patat et al. 2001, Ap.J. 555, 900; SN 1998S at phases +56 and +110 days, Fasia et al. 2001, MNRAS 325, 907) and the 0.924-μm emission as O I 0.9261-μm (e.g., SN 1998S at phase +72.3 days, Fasia et al., op.cit.).

COMETS C/1996 R3, P/2003 A1, AND P/2004 A1 The IAU Committee on Small Bodies Nomenclature has decided to

name three comets as follows: C/1996 R3 (Lagerkvist), P/2003 A1 (LIN-EAR), P/2004 A1 (LONEOS).

2004 November 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green